OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#14793 — men75-1-dsl
Incident Report for Web Cloud
Resolved
The DSLAM men75-1-dsl is no longer responding to ping.

Update(s):

Date: 2015-09-24 13:32:59 UTC
Our teams ate still on site trying to locate the cut.

Date: 2015-09-24 13:32:22 UTC
Ongoing incident in the PARIS MENILMONTANT sector.
We have operators in the field.

Date: 2015-09-24 13:31:21 UTC
The site is isolated.

There is a fiber cut between men75 sites and th2, and between men75 sites and bgl93.

A ticket has been generated by our fiber operator.
Posted Sep 24, 2015 - 13:30 UTC